OpenStudy (anonymous):

Stumped and really need some instruction... Confirm that f and g are inverses by showing that f(g(x))=x and g(f(x))=x f(x)=2/x g(x)=2/x

jaynator495 (jaynator495):

well... you just stumped me O_O sorry XD sorry i wasnt much help =/, -Jaynator495

OpenStudy (anonymous):

This is my work so far... f(g(x))=x g(f(x))=x 2/(2/x) 2/(2/x)

OpenStudy (anonymous):

absolutely

OpenStudy (anonymous):

is there a way to solve this?

OpenStudy (kirbykirby):

Well what you did just needs simplification :)

OpenStudy (kirbykirby):

\[\Large \frac{2}{\frac{2}{x}}=2\div \frac{2}{x}=2\times \frac{x}{2} \]

OpenStudy (anonymous):

so the answer is 2x/2?

OpenStudy (kirbykirby):

(sorry I read x/2 at first hehe)

OpenStudy (anonymous):

so is my answer correct?

OpenStudy (kirbykirby):

yeah \(\checkmark \)

OpenStudy (anonymous):

thanks!